Consider how Pine Valley​, a popular ski​ resort, could use capital budgeting to decide whether the $10.5 million. Spring Park Lodge expansion would be a good investment.

DATA TABLE

Assume that

Pine Valley​'s managers developed the following estimates concerning a planned expansion to its Spring Park Lodge​ (all numbers​ assumed):

Number of additional skiers per day. . . . . . . . . . . . . . . . .

122

Average number of days per year that weather

conditions allow skiing at Pine Valley. . . . . . . . . .

165

Useful life of expansion (in years). . . . . . . . . . . . . . . . . . .

10

Average cash spent by each skier per day. . . . . . . . . . .

$244

Average variable cost of serving each skier per day. . . .

$136

Cost of expansion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

$10,500,000

Discount rate. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

14%

Assume that Pine Valley uses the​ straight-line depreciation method and expects the lodge expansion to have a residual value of $700,000 at the end of its ten​-year life.

1.

Compute the average annual net cash inflow from the expansion.

2.

Compute the average annual operating income from the expansion.

3.

Compute the payback period.

4.

Compute the ARR.
Answer #1

solution 1:

Average annual net cash inflow from operation = Nos of skiers day * Contribution margin per skier

(122*165) * ($244 - $136) = $2,174,040

Solution 2:

Average annual operating income from expansion = Annual cash inflow - Depreciation

= $2,174,040 - ($10,500,000 - $700,000) / 10= $1,194,040

Solution  3:

Payback period = Initial investment / Annual cash inflows = $10,500,000 / $2,174,040 = 4.83 years

Solution 4:

ARR = Average annual income / Average investment

Average investment = (Cost +Residual value) / 2 = ($10,500,000 / $700,000) / 2 = $5,600,000

ARR = $1,194,040 / $5,600,000 = 21.32%